HELPMEASAPPP:

Factorise: 1) 12-(x+x²) (8-x-x²) 2) 9(x-y)²-(x+2y)²

jhonyy9:

do you think factorise completely ?

jhonyy9:

but first of all this is a math question please re-post it on the math subject - ty.

Vocaloid:

Moved to mathematics

HELPMEASAPPP:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid Moved to mathematics \(\color{#0cbb34}{\text{End of Quote}}\) Alr

HELPMEASAPPP:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 do you think factorise completely ? \(\color{#0cbb34}{\text{End of Quote}}\) Wym?

surjithayer:

1. \[12-(x+x^2)[8-(x+x^2)]\] \[=12-8(x+x^2)+(x+x^2)^2\] \[put~x+x^2=y\] \[y^2-8y+12\] \[=y^2-2y-6y+12\] \[=y(y-2)-6(y-2)\] \[=(y-2)(y-6)\] \[=(x^2+x-2)(x^2+x-6)\] \[=[x^2+2x-x-2][x^2+3x-2x+6]\] \[=[x(x+2)-1(x+2)][x(x+3)-2(x+3)]\] \[=(x+2)(x-1)(x+3)(x-2)\] \[=(x-2)(x-1)(x+2)(x+3)\] 2. \[9(x-y)^2-(x+2y)^2\] \[=[3(x-y)]^2-(x+2y)^2\] \[=(3x-3y)^2-(x+2y)^2\] \[=[3x-3y+x+2y][3x-3y-x-2y]\] \[=(4x-y)(2x-5y)\]

HELPMEASAPPP:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @surjithayer 1. \[12-(x+x^2)[8-(x+x^2)]\] \[=12-8(x+x^2)+(x+x^2)^2\] \[put~x+x^2=y\] \[y^2-8y+12\] \[=y^2-2y-6y+12\] \[=y(y-2)-6(y-2)\] \[=(y-2)(y-6)\] \[=(x^2+x-2)(x^2+x-6)\] \[=[x^2+2x-x-2][x^2+3x-2x+6]\] \[=[x(x+2)-1(x+2)][x(x+3)-2(x+3)]\] \[=(x+2)(x-1)(x+3)(x-2)\] \[=(x-2)(x-1)(x+2)(x+3)\] 2. \[9(x-y)^2-(x+2y)^2\] \[=[3(x-y)]^2-(x+2y)^2\] \[=(3x-3y)^2-(x+2y)^2\] \[=[3x-3y+x+2y][3x-3y-x-2y]\] \[=(4x-y)(2x-5y)\] \(\color{#0cbb34}{\text{End of Quote}}\) Thnx

surjithayer:

yw